Circular No. 5749

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRB 930309 The COMPTEL team (K. Bennett, Astrophysics Division, Space Sci-

ence Department, ESTEC, European Space Agency; V. Schonfelder, Max-Planck-Institut für Extraterrestrische Physik; J. Ryan, University of New Hampshire; W. Hermsen, SRON Laboratory for Space Research, Leiden; and their colleagues) reports: “We have imaged a γ-ray burst at > 1 MeV. The event was detected by BATSE on Mar. 9.13053 UT and was of ∼ 20 s duration in the COMPTEL data. The imaging analysis is preliminary, and because of limited counts, the error in position is larger than the nominal COMPTEL resolution of 1^{o}. The most likely position of the burst is α = 21^{h}32^{m}31^{s}, δ = +54^{o}39′.0 (equinox 2000.0). The four corners of the 1σ error box are: 21^{h}40^{m}00^{s}, +55^{o}25′.2; 21^{h}38^{m}07^{s}, +53^{o}15′.6; 21^{h}27^{m}07^{s}, +53^{o}19′.8; 21^{h}27^{m}22^{s}, +55^{o}30′.0. Because of the proximity of the sun in early March, the field was poorly observable with optical instruments. Only radio observatories who expressed interest were notified as part of the COMPTEL/NMSU Rapid Burst Response Campaign. However, new observations are encouraged. Interested observers can contact Ryan (e-mail jryan@unh.edu; telephone 603-862-3510) or B. McNamara at New Mexico State University (e-mail bmcnamar@nmsu.edu; telephone 505-646-2614).”

L. Hanlon and K. Bennett (ESA/ESTEC), with T. Spoelstra (Dwingeloo), communicate: “We report the detection of a possible radio counterpart of the γ-ray burst GRB 930309. The radio source is located at α = 21^{h}32^{m}39^{s}.0, δ = +54^{o}38′34′′ (equinox 2000.0; the positional error is within the least significant digit). This is close to the COMPTEL maximum likelihood position given above. Radio observations of the GRB error box began on Mar. 12 (at 49 cm), with subsequent observations on Mar. 16 (49 cm) and 18 (6 cm). The source was initially observed to have a flux of 76 ± 3.6 mJy at 49 cm. Three days later, the flux had declined to 66 ± 2.9 mJy. The observation at 6 cm gives an upper limit on the flux of 1 mJy, indicating that the source has an unusually steep spectral slope. Monitoring at 49 cm is still in progress. In view of the pressing need to identify the nature of γ-ray bursts we encourage complementary observations of this source at other wavelengths.”

1993 April 5 Daniel W. E. Green